# Retrieve Anyone: A General-purpose Person Re-identification Task with Instructions

### v1

@@ 7 Conclusion and Limitations
In this paper, we tackle existing traditional ReID (Trad-ReID), clothes-changing ReID (CC-ReID), clothes template based clothes-changing ReID (CTCC-ReID) and language-instruct ReID (LI-ReID) by proposing a new instruct-ReID task, which holds great potential and broad impact in the security application, scientific research and other fields. For a specific task case, the target person images in the gallery are retrieved by jointly considering the query person image and given instructions. Based on instruct-ReID, we establish a large-scale and comprehensive OmniReID benchmark with the existing publicly available datasets. Furthermore, we design a generic framework RAM for the proposed instruct-ReID with an adaptive triplet loss to consider identity and instruction alignments. One limitation of this work is that instruct-ReID only implements four ReID tasks with task-specific instructions. We hope our OmniReID can facilitate future works such as unified network structure design and multi-task learning methods on a broad variety of retrieval tasks.
@@
Appendix A Details of Language Annotation Generation
@@
A.1 Pedestrian Attribute Generation
We annotate 20 attributes and 92 specific representation in words for OmniReID, as shown in Table. 5. The attributes are carefully selected considering a wide range of human visual characteristics from the datasets, including full-body clothing, hair color, hairstyle, gender, age, actions, posture, and accessories such as umbrellas or satchels. Though there might be more than one representation for attributes such as coat color and trousers color, only one representation corresponding to the image is selected for annotation. The attributes are annotated by professional annotators in the image level, thus the annotation file will contain more accurate and detailed description. For example, in Fig. 6(a), although the two images are of the same identity, the difference of pedestrian angle is also annotated.
@@ A.1 Pedestrian Attribute Generation
"black coat", "blue coat", "gray coat", "green coat", "purple coat", "red coat", "white coat", "yellow coat"
@@
"black trousers", "blue trousers", "gray trousers", "green trousers", "purple trousers", "red trousers", "white trousers", "yellow trousers"
@@
"agnostic length coat", "long sleeve coat", "short sleeve coat", " bareback coat"
@@
"shorts trousers", "skirt", "trousers"
@@
"female", "agnostic gender", "male"
@@
"without glasses", "with glasses", "with sunglasses"
@@
"black hair", "agnostic color hair", "white hair", "yellow hair"
@@
"bald hair", "agnostic style hair", "long hair", "short hair"
@@
"backpack", "hand bag", "shoulder bag", "waist pack", "trolley", "agnostic style bag", "without bag"
@@
"black shoes", "blue shoes", "gray shoes", "green shoes", "purple shoes", "red shoes", "white shoes", "yellow shoes"
@@
"boots", "leather shoes", "sandal", "walking shoes"
@@
"back", "front", "side"
@@
"lie", "pose agnostic", "sit", "stand", "stoop"
@@
"business suit", "agnostic style coat", "dress", "jacket", "long coat", "shirt", "sweater", "t-shirt"
@@
"with glove", "agnostic glove", "without glove"
@@
"smoking", "agnostic smoking", "without smoking"
@@
"with umbrella", "without umbrella"
@@
"chef uniform", "common clothing", "firefighter uniform", "medical uniform", "office uniform", "agnostic uniform", "worker uniform"
A.2 Attribute-to-Language Transformation
We provide some example from our OmniReID that provides sentence descriptions of individuals in images. Compared with discrete attribute words, language is more natural for consumers. To this end, we transform these attributes into multiple sentences using the Alpaca-LoRA large language model. Specifically, we ask the Alpaca-LoRA with the following sentences: "Generate sentences to describe a person. The above sentences should contain all the attribute information I gave you in the following." Annotators carefully check the generated annotations to ensure the correctness of the language instructions. Fig. 6a presents the examples of the same identity and Fig. 6b presents the examples of the transformation with different domains and identities.
Figure 6: We first generate attributes for a person and then transform attributes into sentences by a large language model. (a) The attributes and language guidance of same identity. (b) More instances of attributes-to-language transformation with different domains and identities.
Appendix B The Discriminative Abilities of FashionCLIP and CLIP
In this section, we compare the capacity of fixed FashionCLIP model [5] and CLIP model [33] as instruction encoder to distinguish different instructions. As shown in Fig. 7, the instruction features are extracted from upper clothes images and we calculate the similarity between template upper clothes and cropped upper clothes images. CLIP model is pretrained on classification dataset which is lack of discernment of human details. Take the first row of Template1 for instance, the similarity score between template clothes and corresponding cropped clothes (in green box) is only 0.698, which is even smaller than the second, forth, fifth and sixth cropped clothes images. The counter-intuitive result indicates that the CLIP model is unable to distinguish texture, color and other style characteristics. The FashionCLIP is a CLIP-like model for the fashion industry and fine-tuned on fashion retrieval dataset. The similarity scores of each template between template clothes and corresponding cropped clothes in the second row are obviously exceed others, which indicates that FashionCLIP model [5] shows better discriminative abilities than CLIP model [33].
Figure 7: We extract instruction features respectively by CLIP model or FashionCLIP model and calculate the instruction similarity scores of upper clothes templates and the cropped upper clothes. Green boxes mean the template clothes images is searched on the internet by these corresponding cropped clothes images.
Appendix C Instruction Generation
@@
C.1 Clothes Template Based Clothes-Changing ReID
@@
C.2 Language-instructed ReID
We provide more examples for LI-ReID as shown in Fig. 8. Similar to CTCC-ReID, the instruction of gallery image is several sentences describing pedestrian attributes. We randomly select the description languages from the corresponding person images in gallery and provide to query images as instruction. The model is required to retrieve images of the same person following the provided sentences.
@@ C.2 Language-instructed ReID
Appendix D Details of OmniReID
@@
D.1 Dataset Statistics of OmniReID
OmniReID collects 8 publicly available datasets of 4 existing ReID tasks, including Traditional ReID (Trad-ReID), Clothes-Changing ReID (CC-ReID), Clothes Template Based Clothes-Changing ReID (CTCC-ReID) and Language-Instructed ReID (LI-ReID). As shown in Table 6, we perform two training scenarios based on the built benchmark OmniReID: 1) Single-task Learning (STL): Every dataset is treated as a single task, which is trained and tested individually. 2) Multi-task Learning (MTL): The model is optimized by joint training of all the ReID tasks with all the training datasets. The trained model is then evaluated on different tasks with various datasets. For the Trad-ReID, we select widely-used MSMT17, CUHK03 and Market1501 datasets. For CC-ReID Task, we select widely-used PRCC, LTCC and VC-Clothes, considering the variety of domains. For CTCC-ReID and LI-ReID tasks, we label the COCAS+ Real1, COCAS+ Real2 datasets with language instructions and clothes image instructions for training and test. OmniReID forms a total of 121786 images for training and 166055 images for test, which unites all-purpose person retrieval into one instruct-ReID task.
Table 6: Dataset statistics of OmniReID. We summarize both the training and test set images based on the built benchmark. $\star$ means the images is the same as the LI-ReID task.
Multi-task Single-task Training Test
@@ D.1 Dataset Statistics of OmniReID
Trad-ReID(MTL) Trad-ReID(STL) MSMT17 [40] 30248 MSMT17 [40] 93820
Trad-ReID(STL) CUHK03 [25] 7365 CUHK03 [25] 6732
Trad-ReID(STL) Market1501 [48] 12936 Market1501 [48] 23100
LI-ReID(STL) LI-ReID(STL) COCAS+ Real1 [23] 34469 COCAS+ Real2 [23] 14449
CC-ReID(MTL) CC-ReID(STL) PRCC [43] 17896 PRCC [43] 10800
CC-ReID(STL) VC-Clothes [39] 9449 VC-Clothes [39] 9611
CC-ReID(STL) LTCC [32] 9423 LTCC [32] 7543
CTCC-ReID(MTL) CTCC-ReID(STL) COCAS+ Real1 [23] $34469^{\star}$ COCAS+ Real2 [23] ${}^{\star}$
Total - - 121786 - 166055
D.2 Discussion of Ethical Issues

### v4

@@ 6 Conclusion
This proposes one unified instruct-ReID task to jointly tackle existing traditional ReID, clothes-changing ReID, clothes template based clothes-changing ReID, language-instruct ReID, visual-infrared ReID, and text-to-image ReID tasks, which holds great potential in social surveillance. To tackle the instruct-ReID task, we build a large-scale and comprehensive OmniReID benchmark and a generic framework with an adaptive triplet loss. We hope our OmniReID can facilitate future works such as unified network structure design and multi-task learning methods on a broad variety of retrieval tasks.
@@
Table 6: More results of IRM on different pre-trained models. We train and test using the default image resolution of 256$\times$128 (256), with (384) indicating an image resolution of 384$\times$192. † denotes that the test mode is VIS-to-IR and ‡ denotes IR-to-VIS mode on LLCM.
IRM CTCC-ReID LI-ReID T2I-ReID VI-ReID CC-ReID Trad-ReID
Real2 Real2 CUHK. LLCM† LLCM‡ LTCC PRCC VC-Clo. Market1501 MSMT17 CUHK03
SOTA - 30.0 [28] 14.9 63.9 [3] 65.8 [61] 62.9 [61] 40.8 [12] 45.9 [16] 71.8 [14] 93.0 [69] 71.8 [69] 77.7 [44]
DeiT [45] STL(256) 31.5 30.1 64.2 65.1 62.3 46.3 43.1 82.1 87.9 67.9 77.5
MTL(256) 40.2 38.7 65.7 66.2 62.9 53.2 46.8 76.3 90.0 68.9 78.7
ALBEF [22] STL(256) 30.6 28.7 67.7 61.2 58.8 40.2 42.1 67.1 78.9 60.3 69.2
MTL(256) 37.2 34.5 68.3 62.1 60.1 41.2 44.1 60.5 79.7 62.1 69.9
HAP [58] STL(256) 30.8 31.1 63.2 66.2 62.9 45.3 44.7 79.2 90.2 70.3 79.6
MTL(256) 39.2 39.8 65.1 67.1 64.3 49.2 48.7 77.3 91.7 72.3 82.2
PASS [69] STL(256) 32.2 30.7 65.3 66.6 64.5 46.7 46.0 80.1 92.3 71.9 83.3
MTL(256) 41.7 39.8 66.5 68.5 67.2 52.0 52.3 78.9 93.5 72.4 85.4
STL(384) 33.7 33.8 65.9 68.9 64.7 48.9 49.3 82.5 92.7 73.7 84.1
MTL(384) 42.5 42.3 67.2 71.2 67.3 54.1 58.4 81.8 93.9 75.4 88.7
Appendix A More results of IRM on different pre-trained models
In this section, we provide more results about IRM on the publicly released pre-trained models i.e., DeiT [45], ALBEF [22], HAP [58], and PASS [69] in Tab. 6. We can see two conclusions from Tab. 6. First, we can see models pre-trained on human-centric images, i.e., HAP [58] and PASS [69], can naturally increase the performance of person-retrieval tasks. Second, although ALBEF [22] exhibits lower performance on Trad-ReID and other image-based ReID tasks, it achieves the highest performance on the T2I-ReID task due to its vision-language pretraining knowledge. Additionally, we validate the effectiveness of IRM on 384$\times$192 (384) image resolution using the PASS pre-trained model. The results show a further improvement in performance compared to the 256$\times$128 results, which indicates that increasing the image size can contribute to achieving better retrieval results.
Appendix B Visualizaton
We visualize the retrieval results of CTCC-ReID, LI-ReID, CC-ReID and Trad-ReID tasks in Fig. 5 and VI-ReID, T2I-ReID in Fig. 6. Given a query image, IRM not only retrieves the right person from the gallery but also finds specific target images of the person following the instruction. Concretely, for CTCC-ReID, IRM retrieves images of query persons wearing instructed clothes as shown in the first row. For LI-ReID, IRM effectively parses information from languages such as bag condition (e.g., row 2 person #1) and clothes attribute (e.g., row 2 person #2,3) to find the correct image. For CC-ReID, with the “Ignore clothes” instruction, our method focuses on biometric features and successfully retrieves the person in the case of changing clothes, e.g. the 4th and 5th images of row 3 person #1, which the clothes are different from the query image. For Trad-ReID shown in row 4, “do not change clothes” instructs IRM to pay attention to clothes, a main feature of a person’s image. In this case, IRM retrieves images with the same clothes in the query image. For VI-ReID, we visualize the retrieval results for both VIS-to-IR and IR-to-VIS modes. IRM can retrieve the correct cross-modality images even in low visual conditions. For T2I-ReID, when there are images of different identities in the gallery but with representations consistent with the text description (e.g., row 1 3rd and row 2 person 4th images in Fig 6(c)), it may introduce some noise to IRM. However, IRM is still capable of indexing most of the correct results based on the given descriptions.
Figure 5: Illustration of all tasks retrieval results. We visualize the task-specific instructions on three people as examples. Green and red boxes mean true and false matches.
Figure 6: Illustration of VI-ReID and T2I-ReID tasks retrieval results. We visualize both VIS-to-IR and IR-to-VIS mode results on three people in the LLCM dataset. There are about four images for each person in CUHK-PEDES, we visualize the top 4 results for T2I-ReID on the CUHK-PEDES dataset as examples. Green and red boxes mean true and false matches.
Appendix C Details of Language Annotation Generation
@@
C.1 Pedestrian Attribute Generation
We annotate 20 attributes and 92 specific representation in words for OmniReID, as shown in Tab. 7. The attributes are carefully selected considering a wide range of human visual characteristics from the datasets, including full-body clothing, hair color, hairstyle, gender, age, actions, posture, and accessories such as umbrellas or satchels. Though there might be more than one representation for attributes such as coat color and trousers color, only one representation corresponding to the image is selected for annotation. The attributes are annotated by professional annotators in the image level, thus the annotation file will contain more accurate and detailed description. For example, in Fig. 7 (a), although the two images are of the same identity, the difference of pedestrian angle is also annotated.
@@ C.1 Pedestrian Attribute Generation
”black coat”, ”blue coat”, ”gray coat”, ”green coat”, ”purple coat”, ”red coat”, ”white coat”, ”yellow coat”
@@
”black trousers”, ”blue trousers”, ”gray trousers”, ”green trousers”, ”purple trousers”, ”red trousers”, ”white trousers”, ”yellow trousers”
@@
”agnostic length coat”, ”long sleeve coat”, ”short sleeve coat”, ” bareback coat”
@@
”shorts trousers”, ”skirt”, ”trousers”
@@
”female”, ”agnostic gender”, ”male”
@@
”without glasses”, ”with glasses”, ”with sunglasses”
@@
”black hair”, ”agnostic color hair”, ”white hair”, ”yellow hair”
@@
”bald hair”, ”agnostic style hair”, ”long hair”, ”short hair”
@@
”backpack”, ”hand bag”, ”shoulder bag”, ”waist pack”, ”trolley”, ”agnostic style bag”, ”without bag”
@@
”black shoes”, ”blue shoes”, ”gray shoes”, ”green shoes”, ”purple shoes”, ”red shoes”, ”white shoes”, ”yellow shoes”
@@
”boots”, ”leather shoes”, ”sandal”, ”walking shoes”
@@
”back”, ”front”, ”side”
@@
”lie”, ”pose agnostic”, ”sit”, ”stand”, ”stoop”
@@
”business suit”, ”agnostic style coat”, ”dress”, ”jacket”, ”long coat”, ”shirt”, ”sweater”, ”t-shirt”
@@
”with glove”, ”agnostic glove”, ”without glove”
@@
”smoking”, ”agnostic smoking”, ”without smoking”
@@
”with umbrella”, ”without umbrella”
@@
”chef uniform”, ”common clothing”, ”firefighter uniform”, ”medical uniform”, ”office uniform”, ”agnostic uniform”, ”worker uniform”
C.2 Attribute-to-Language Transformation
We provide some example from our OmniReID that provides sentence descriptions of individuals in images. Compared with discrete attribute words, language is more natural for consumers. To this end, we transform these attributes into multiple sentences using the Alpaca-LoRA large language model. Specifically, we ask the Alpaca-LoRA with the following sentences: ”Generate sentences to describe a person. The above sentences should contain all the attribute information I gave you in the following.” Annotators carefully check the generated annotations to ensure the correctness of the language instructions. Fig. 7 (a) presents the examples of the same identity and Fig. 7 (b) presents the examples of the transformation with different domains and identities.
Figure 7: We first generate attributes for a person and then transform attributes into sentences by a large language model. (a) The attributes and language guidance of the same identity. (b) More instances of attributes-to-language transformation with different domains and identities.
Appendix D Instruction Generation
@@
D.1 Instructions on traditional ReID, cloth-changing ReID and visual-infrared ReID
Following the methods in Instruct-BLIP [8], which uses GPT-4 to generate 20 different instructions for traditional ReID (e.g., We use ’do not change clothes’ to generate equivalent expressions such as ’maintain consistent clothes’ and so on), 20 different instructions for clothes-changing ReID (e.g., We use ’ignore clothes’ to generate equivalent expressions such as ’change your clothes’ and so on), and 20 different instructions for visual-infrared ReID (e.g., We use ’retrieve cross-modality images’ to generate equivalent expressions such as ’fetch images across different modalities’ and so on) as shown in Tab. 8 in detail. We randomly choose one from these instructions when training each mini-batch and evaluating every instruction for testing the model.
Table 8: Details of instructions for Trad-ReID, CC-ReID, and VI-ReID in OmniReID. We use GPT-4 to generate 20 different expressions for ’do not change clothes‘ as instructions for traditional ReID and similarly, we generated 20 different expressions of ’ignore clothes’ for clothes-changing ReID and 20 different expressions of ’retrieve cross-modality images’ for visual-infrared ReID.
ReID task instruction representation
Trad-ReID
”do not change clothes”, ”maintain consistent clothes”, ”keep original clothes”, ”preserve current clothes”, ”retain existing clothes”, ”wear the same clothes”, ”stick with your clothes”, ”don’t alter your clothes”, ”no changes to clothes”, ”unchanged outfit”, ”clothes remain constant”, ”no clothing adjustments”, ”steady clothing choice”, ”clothing remains unchanged”, ”consistent clothing selection”, ”retain your clothing style”, ”clothing choice remains”, ”don’t swap clothes”, ”maintain clothing selection”, ”clothes stay the same”
CC-ReID
”change your clothes”, ”swap outfits”, ”switch attire”, ”get into a different outfit”, ”try on something new”, ”put on fresh clothing”, ”dress in alternative attire”, ”alter your outfit”, ”wear something else”, ”don a different ensemble”, ”trade your garments”, ”shift your wardrobe”, ”exchange your clothing”, ”update your attire”, ”replace your outfit”, ”clothe yourself differently”, ”switch your style”, ”update your look”, ”put on a new wardrobe”, ”ignore clothes”
VI-ReID
”retrieve cross-modality images”, ”fetch images across different modalities”, ”collect images from various modalities”, ”obtain images spanning different modalities”, ”retrieve images from diverse modalities”, ”gather images across modalities”, ”access images across different modalities”, ”acquire images spanning various modalities”, ”extract images from different modalities”, ”retrieve images across multiple modalities”, ”fetch images from distinct modalities”, ”collect images across various modalities”, ”access images from different modalities”, ”obtain images from diverse modalities”, ”gather images spanning different modalities”, ”extract images from various modalities”, ”retrieve images across varied modalities”, ”obtain images from distinct modalities”, ”access images across multiple modalities”, ”collect images spanning diverse modalities”
D.2 Text to image ReID
In T2I-ReID scenario, during the training process, both images and responding description texts are fed into IRM. We adopt a contrastive loss to align the image features and text features. To further enhance the retrieval capability of the model, we employ a classifier to determine whether an inputted image-text pair is positive or negative. Specifically, we retain the original image-text pairs as positives and form negative pairs by matching text features with unrelated image features before inputting them into the attention module. In the inference stage, the query is the describing sentences, and the image features and query features are extracted separately. given a query text feature, we rank all the test gallery image features based on their similarity with the text. We select the top 128 image features and pair them with the query text feature. These pairs are then input into the attention module, further utilizing the matching scores to rank these images. The search is deemed to be successful if top-K images contain any corresponding identity.
D.3 Clothes Template Based Clothes-Changing ReID
@@
D.4 Language-instructed ReID
We provide more examples for LI-ReID as shown in Fig. 9. Similar to CTCC-ReID, the instruction for gallery images is several sentences describing pedestrian attributes. We randomly select the description languages from the corresponding person images in gallery and provide to query images as instruction. The model is required to retrieve images of the same person following the provided sentences.
@@ D.4 Language-instructed ReID
Appendix E Details of OmniReID
@@
Table 9: Dataset statistics of OmniReID. We summarize both the training and test set images based on the built benchmark. †means the images is the same as the LI-ReID task, we do not calculate the same image twice when calculating the total number of images in Multi-task Learning. ‡means the LTCC with extra clothes templates for CTCC task.
Single-task learning Training Test
@@
Trad-ReID MSMT17 [47] 30,248 MSMT17 [47] 93,820
Trad-ReID CUHK03 [30] 7,365 CUHK03 [30] 6,732
Trad-ReID Market1501 [64] 12,936 Market1501 [64] 23,100
LI-ReID COCAS+ Real1 [28] 34,469 COCAS+ Real2 [28] 14,449
CC-ReID PRCC [52] 17,896 PRCC [52] 10,800
CC-ReID VC-Clothes [46] 9,449 VC-Clothes [46] 9,611
CC-ReID LTCC [37] 9,423 LTCC [37] 7,543
CTCC-ReID COCAS+ Real1 [28] 34,469† COCAS+ Real2 [28] 14,449†
VI-ReID LLCM [61] 30,921 LLCM [61] 13,909
T2I-ReID CUHK-PEDES [26] 28,566 CUHK-PEDES [26] 3,074
Multi-task learning Training Test
Dataset Number of images Dataset Number of images
OmniReID MSMT17 [47] 4,973,044 MSMT17 [47] 93,820
+CUHK03 [30] CUHK03 [30] 6,732
+Market1501 [64] Market1501 [64] 23,100
+COCAS+ Real1 [28] COCAS+ Real2 [28] 14,449
+PRCC [52] PRCC [52] 10,800
+VC-Clothes [46] VC-Clothes [46] 9,611
+LTCC [37] LTCC [37] 7,543
+COCAS+ Real1 [28] COCAS+ Real2 [28] 14,449†
+LLCM [61] LLCM [61] 13,909
+CUHK-PEDES [26] CUHK-PEDES [26] 3,074
+LTCC‡ [37] - -
+SYNTH-PEDES [70] - -
E.1 Dataset Statistics of OmniReID
OmniReID collects 12 publicly available datasets of 6 existing ReID tasks, including Traditional ReID (Trad-ReID), Clothes-Changing ReID (CC-ReID), Clothes Template Based Clothes-Changing ReID (CTCC-ReID), Visual-Infrared ReID (VI-ReID), Text-to-Image ReID (T2I-ReID) and Language-Instructed ReID (LI-ReID). As shown in Tab. 9, we perform two training scenarios based on the built benchmark OmniReID: 1) Single-task Learning (STL): Every dataset is treated as a single task, which is trained and tested individually. 2) Multi-task Learning (MTL): The model is optimized by joint training of all the ReID tasks with all the training datasets. The trained model is then evaluated on different tasks with various datasets.In Trad-ReID, we utilize the widely-used MSMT17, CUHK03, and Market1501 datasets. For the CC-ReID task, we choose the widely-used PRCC, LTCC, and VC-Clothes datasets, considering the diversity of domains. Regarding the CTCC-ReID task, we label LTCC with clothes image instructions and employ COCAS+ Real1, LTCC for training, while COCAS+ Real2 serves as the test set. For LI-ReID, we annotate the COCAS+ Real1 and COCAS+ Real2 datasets with language instructions for training and test. In VI-ReID, we select a new and challenging low-light cross-modality dataset called LLCM. Finally, for T2I-ReID, we opt for the widely-used CUHK-PEDES and SYNTH-PEDES datasets for training, with CUHK-PEDES serving as the test set. OmniReID forms a total of 4973044 images for training and 183038 images for test, which unites all-purpose person retrieval into one instruct-ReID task.
OmniReID Agreement This Agreement outlines the terms and conditions governing the use of OmniReID. By signing this agreement, the Recipient agrees to the following terms: The Recipient agrees the OmniReID does not claim Copyrights of Market1501, CUHK03, MSMT17, PRCC, VC-Clothes, LTCC, COCAS+, LLCM, CUHK-PEDES, SYNTH-PEDES and they SHOULD obtain these public datasets from these data providers. The Recipient agrees they must comply with ALL LICENSEs of Market1501, CUHK03, MSMT17, PRCC, VC-Clothes, LTCC, COCAS+, LLCM, CUHK-PEDES, SYNTH-PEDES when they use OmniReID. The Recipient agrees the demographic makeup of OmniReID is not representative of the broader population. The Recipient agrees OmniReID should only be available for non-commercial research purposes. Any other use, in particular any use for commercial purposes, is prohibited. The Recipient agrees not to use the data for any unlawful, unethical, or malicious purposes. The Recipient agrees not to further copy, publish or distribute any portion of the OmniReID. The Recipient agrees [OUR INSTITUTE] reserves the right to terminate the access to the OmniReID at any time. Name:
Organization/Affiliation:
Position:
Email:
Address:
Address (Line2):
City: Country:
Signature:
Date:
E.2 Discussion of Ethical Issues